Find the amount of interest earned for a deposit of $24,677.23 at 9.72% compounded monthly for 32 months.
Answer by rfer(16322)

A=24677.23(1+0.0972/12)^32
A=24677.23(1.0081)^32
A=24677.23*1.29454
A=31945.67
-------------
31945.67-24677.23=$7,268.44 interest earned